# AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications

### v2

@@ V-B1 Closing real-to-sim-to-real gap
A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments. Models trained in simulators might likely experience performance degradation when tested in the real world, where the environment includes unpredictable variations that simulators cannot fully replicate. Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training. What distinguishes an DT from a conventional computer simulator? We believe the key lies in its capability of generating simulations and predictions consistent with the real world, as well as preserving intervention performance. Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world). The smaller these two gaps are, the closer an DT is to reality.
@@
Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [222] and policy learning [223, 224, 225, 226]. The major application area of sim2real transfer in DT is robotic manipulation [227], while that in urban navigation and autonomous driving has witnessed a gradual surge, participially in computer vision [228] and reinforcement learning [229, 230]. Key methods include domain randomization [231, 207], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [232, 233], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
@@
How do we characterize real2sim gap and control such a gap? This boils down to quantifying errors of the digital representation of a physical world. Minimizing the real2sim gap is key to system identification and representation learning. Depending on observability and internal workings, a system could be modeled as white-box, grey-box, or black-box [234]. Domain knowledge could help represent the physical world with higher accuracy. For example, hybrid twin [235, 236] relies on both data-driven and physics-informed AI [237]. Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular. Digital shadow [11], a digital model with one-way data exchange from the physical to the digital, takes less effort to build, but could fail to update the state of the physical world once feedback is executed. Digital cousin [238] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
V-B2 Prototyping DTs for human intelligence

### v3

@@ 6.2.1 Closing real-to-sim-to-real gap
A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments. Models trained in simulators might likely experience performance degradation when tested in the real world, where the environment includes unpredictable variations that simulators cannot fully replicate. Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training. Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world). The smaller these two gaps are, the closer an DT is to reality.
@@
Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [165] and policy learning [164, 178, 134, 36]. The major application area of sim2real transfer in DT is robotic manipulation [1], while that in urban navigation and autonomous driving has witnessed a gradual surge, particularly in computer vision [229] and reinforcement learning [83, 109]. Key methods include domain randomization [193, 147], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [195, 227], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
@@
How do we characterize real2sim gap and control such a gap? This boils down to quantifying errors of the digital representation of a physical world. Minimizing the real2sim gap is key to system identification and representation learning. Depending on observability and internal workings, a system could be modeled as white-box, grey-box, or black-box [98]. Domain knowledge could help represent the physical world with higher accuracy. For example, hybrid twin [33] relies on both data-driven and physics-informed AI. Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular. Digital shadow [74], a digital model with one-way data exchange from the physical to the digital, takes less effort to build, but could fail to update the state of the physical world once feedback is executed. Digital cousin [37] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
6.2.2 Prototyping DTs for human intelligence
@@ 6.2.3 Limits in AI
Acknowledgments
This work was supported by NSF CPS-2038984 and ERC-2133516. We thank Mengxuan Liu for her assistance in generating the 3D models.